# Exponentially Enhanced Scheme for the Heralded Qudit GHZ State in Linear Optics

### v1

@@ 3. $X_{d}$: The $d$-level unitary generalization of $\sigma_{x}$ in qubit,
$\frac{1}{3^{9}}\Big{[}$ $6^{3}{\hat{a}^{\dagger}}_{0\tilde{0}}{\hat{a}^{\dagger}}_{2\tilde{1}}{\hat{a}^{\dagger}}_{1\tilde{2}}{\hat{b}^{\dagger}}_{0\tilde{0}}{\hat{b}^{\dagger}}_{2\tilde{1}}{\hat{b}^{\dagger}}_{1\tilde{2}}{\hat{c}^{\dagger}}_{0\tilde{0}}{\hat{c}^{\dagger}}_{2\tilde{1}}{\hat{c}^{\dagger}}_{1\tilde{2}}$
$+6^{3}{\hat{a}^{\dagger}}_{2\tilde{0}}{\hat{a}^{\dagger}}_{1\tilde{1}}{\hat{a}^{\dagger}}_{0\tilde{2}}{\hat{b}}^{\dagger}_{2\tilde{0}}{\hat{b}}^{\dagger}_{1\tilde{1}}{\hat{b}}^{\dagger}_{0\tilde{2}}{\hat{c}^{\dagger}}_{2\tilde{0}}{\hat{c}^{\dagger}}_{1\tilde{1}}{\hat{c}^{\dagger}}_{0\tilde{2}}$
$-(3\sqrt{3})^{3}\alpha^{3}\hat{X}_{\tilde{1}}\hat{Y}_{\tilde{1}}\hat{Z}_{\tilde{1}}({\hat{a}^{\dagger}}_{1,\tilde{1}}{\hat{a}^{\dagger}}_{2,\tilde{1}}+{\hat{a}^{\dagger}}_{1,\tilde{2}}{\hat{a}^{\dagger}}_{2,\tilde{0}})$
×$~{}~{}~{}\times({\hat{b}^{\dagger}}_{1,\tilde{1}}{\hat{b}^{\dagger}}_{2,\tilde{1}}+{\hat{b}^{\dagger}}_{1,\tilde{2}}{\hat{b}^{\dagger}}_{2,\tilde{0}})({\hat{c}^{\dagger}}_{1,\tilde{1}}{\hat{c}^{\dagger}}_{2,\tilde{1}}+{\hat{c}^{\dagger}}_{1,\tilde{2}}{\hat{c}^{\dagger}}_{2,\tilde{0}})\Big{]}|vac\rangle$ (12)
After the postselection with $D_{F}$s and feedforwards, the final state becomes
$\frac{6^{3}}{3^{9}}\Big{[}|0,0,0\rangle-3\sqrt{3}\alpha^{3}|1,1,1\rangle+|2,2,2\rangle\Big{]}.$ (13)
Setting $\alpha=\frac{1}{\sqrt{3}}$, the final state becomes the tripartite qutrit GHZ state. The success probability is given by
$P_{suc}(3,3)=(\frac{2}{9})^{6}\sim 3.6\times 10^{-4}.$ (14)
Comparison with the linear optical schemes in Ref. [25].—
Compared to the schemes using the zero-transmission law (ZTL) in Fourier transform interferometers in Ref. [25] (which we call ‘ZTL scheme’ from now on), our scheme is substantially more efficient for $N\geq 3$ both in the photon numbers and success probabilities. For $N=2$ (qudit Bell state), our scheme has lower success probability ($d\Big{(}\frac{d!}{d^{d}}\Big{)}^{4}$ is smaller than $\frac{d(2d-1)!}{(2d+1)^{2d-1}}$ of the ZTL scheme) but requires less photons (we need $2d$ photons while the ZTL scheme needs $12d+$ 1).
Since the general ZTL scheme in Ref. [25] has a very low efficiency, an alternative near optimal scheme is proposed that can be applied to general $N$ and $d$. We can see that the near-optimal solution still needs significantly more photons and exponentially lower success probability than ours. Fig. 4 compares the required photon numbers and $P_{suc}$ of the near-optimal ZTL scheme with ours for $N=3$ and $2\leq d\leq 8$. On the other hand, a brute-fore method is also used in Ref. [25] for $N=d=3$ that diminishes the photon number to $25$ with $P_{suc}\sim 10^{-10}$, which is still much less efficient than our scheme that requires $9$ photons with $P_{suc}\sim 3.6\times 10^{-4}$. The authors increased $P_{suc}$ to $\sim 0.8\times 10^{-4}$ with multiplexing, which however is still lower than a quarter of ours.
Figure 4: Comparison of our scheme with the near-optimal scheme of Ref. [25] for $N=3$ in (a) photon numbers, and (b) the logarithmic scales of the success probabilities. Graph (b) shows that the success probability is enhance exponentially in our scheme.
@@
The linear optical logic gates in Fig. 1 can be implemented with any qudit encoding system of linear optics. Among them, we discuss the OAM and multi-rail encodings here.
In the OAM encoding, it is straightforward to establish the scheme, because the the OAM beam splitter [30] and OAM-only Fourier transformation operators [31] implement the $CN_{d}$ and $F_{d}$ gates respectively.
In the multi-rail encoding, a permutation of photon paths becomes the $CN_{d}$ gate and the $d$-partite port that Fourier transforms the spatial states of photons becomes the $F_{d}$ gate and Indeed, for $d=3$ case, the multi-rail implementation of linear optical elements are given by
[…]
which can be directly generalized to any $d$-level systems. Replacing all the linear optical logic gates in Fig. 3 with the operators in (15), we obtain a circuit for the $N=3$ qutrit GHZ state in the multirail encoding as in Fig. 5.
Figure 5: $N=d=3$ GHZ generation circuit in the multirail encoding example.
An advantage of the multirail encoding circuit is that $CN_{d}$ and $\overline{CN}_{d}$ is achieved with the permutation of paths, hence the final circuit consists of simpler optical elements, i.e., $d$-partite ports including BSs, permutations of paths, and photon detectors.
@@
Our scheme has several manifest advantages in using the qudit GHZ state as a resource for high-dimensional quantum information processing. First, as we have already demonstrated, our scheme enables more efficient generations of the resource states with single-photons and linear operators, which becomes a more feasible resource state that can generate qudit cluster states with fusion gates for the qudit MBQC [23, 24, 25]. Second, our scheme is adequate to verify the optimal state for the violation of Bell-CHSH nonlocality test. While the maximally entangled state entails the maximal Bell-CHSH violation in $d=2$ systems, the scenario becomes more complicated for higher dimensional systems. Ref. [12, 13] shows that the maximal Bell-CHSH violation occurs for a non-maximally entangled state, which is also verified numerically for the $N=d=3$ case in Ref. [32]. Therefore, for the experiments of those maximal violations, schemes for generalized qudit GHZ states with relatively asymmetric amplitudes are indispensable. Our scheme easily achieve the condition by varying the transitivity of beamspliters. Third, our scheme consists of linear optical logic gates that any type of qubit encoding (e.g., optical angular momentum (OAM), multi-rail, time-bin, etc.) can realize.
In addition, the methodology of composing linear optical logic gates based on the LQG picture can be useful to find more general types of qudit multipartite entanglement. For example, Ref. [33] proposed heralded schemes for $N$-partite $N$-level symmetric and anti-symmetric states based on the method. We expect to construct more rigorous and general formalism of the LQG picture that will provide a more systematic path for the heralded generations of essential qudit entanglement such as absolutely maximally entangled (AME) states.
@@
SC is grateful to Prof. Jung-Hoon Chun for his support on this research. SC is supported by National Research Foundation of Korea (NRF, 2019R1I1A1A01059964, RS-2023-00245747). JR acknowledges support from the National Research Foundation (NRF) of Korea grant funded by the Korea Government (Grant No. NRF-2022M3K2A1083890). YSK acknowledges support from the KIST institutional program (2E32941).
@@ Appendix A Operations of $CN_{d}$, the $d$-level CNOT gate between the spatial mode and the internal state
Unlike the convential generalized CNOT gate that works on two different qudits, $CN_{d}$ and $\overline{CN}_{d}$ here transform the spatial mode of a given photon according to its internal state. They transform the two variables in the same way as the generalized CNOT gate between two qudits. The spatial mode and internal state play the role of the target and control states.
More explicitly, the operation of $CN_{d}$ given by Eqs. (Linear optical logic gates for generating $|GHZ_{N}^{d}\rangle$.—) can be enumerated as the following table:
$|\tilde{0}\rangle_{0}\to|\tilde{0}\rangle_{0}$ $|\tilde{1}\rangle_{0}\to|\tilde{1}\rangle_{1}$cdots $|\widetilde{d-1}\rangle_{0}\to|\widetilde{d-1}\rangle_{d-1}$
$|\tilde{0}\rangle_{1}\to|\tilde{1}\rangle_{1}$ $|\tilde{1}\rangle_{1}\to|\tilde{1}\rangle_{2}$cdots $|\widetilde{d-1}\rangle_{1}\to|\widetilde{d-1}\rangle_{0}$
@@
$|\tilde{0}\rangle_{d-2}\to|\tilde{0}\rangle_{d-2}$ $|\tilde{1}\rangle_{d-2}\to|\tilde{1}\rangle_{d-1}$cdots $|\widetilde{d-1}\rangle_{d-2}\to|\widetilde{d-1}\rangle_{d-3}$
$|\tilde{0}\rangle_{d-1}\to|\tilde{0}\rangle_{d-1}$ $|\tilde{1}\rangle_{d-1}\to|\tilde{1}\rangle_{0}$cdots $|\widetilde{d-1}\rangle_{d-1}\to|\widetilde{d-1}\rangle_{d-2}$
@@ Appendix B Comprehensive analysis on the qudit $N$-partite GHZ state generation scheme
Here we provide a comprehensive explanation on the linear optical circuit for generating the qudit GHZ state given in Fig. 2. We first explain how we designed our scheme based on the graph picture of linear quantum networks (the LQG picture) [34, 28] and then show that the scheme indeed generates the expected target state.
B.1 Construction of linear optical circuit from the boson subtraction operator
Our scheme can be considered a linear optical realization of the boson subtraction scheme in Ref. [28], which exploited the LQG picture [34, 28] to find boson subtraction schemes for generating multipartite entanglement of qubits and qudits. Those boson subtaction schemes are in principle realized as linear optical heralded circuits, which is elucidated as a set of direct translation rule for qubit cases in Ref. [35]. However, its general extension to arbitrary $d$-level cases is absent.
In Ref. [28], the bipartite graph (bigraph) that corresponds to the boson subtraction scheme for the qudit GHZ state is given by
[…]
which corresponds to the following operator (see Table I of Ref. [28]):
$\hat{A}_{N,d}=\Big{(}\frac{1}{\sqrt{2}}\Big{)}^{(d-1)N}({\hat{a}}_{1,{\tilde{0}}}-{\hat{a}}_{2,\widetilde{d-1}})^{d-1}({\hat{a}}_{2,{\tilde{0}}}-{\hat{a}}_{3,\widetilde{d-1}})^{d-1}({\hat{a}}_{3,{\tilde{0}}}-{\hat{a}}_{4,\widetilde{d-1}})^{d-1}\cdots\times({\hat{a}}_{N,{\tilde{0}}}-{\hat{a}}_{1,\widetilde{d-1}})^{d-1}.$ (16)
We can intuitively see that the following translation holds from the LQG picture to the linear optical system:
@@ B.1 Construction of linear optical circuit from the boson subtraction operator
Therefore, by combing the $N$ sets of the above linear optical element with the structure of the bigraph (B.1), we obtain the scheme in Fig. 2 that generates the expected target state.
@@
B.2 Demonstrations of the final state generated by the scheme in Fig. 2
@@
To demonstrate that the scheme in Fig. 2 generates the qudit GHZ state, we need the following two identities that were first given in Ref. [28] without proofs. Due to their significance in calculating our circuit operation, we restate them here with proofs.
@@
$({\hat{a}}_{{\tilde{0}}})^{l}({\hat{a}}_{\widetilde{d-1}})^{d-1-l}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle=(-1)^{d-1-l}\frac{l!(d-1-l)!}{\sqrt{d}^{d-2}}{\hat{a}}^{\dagger}_{\widetilde{d-1-l}}|vac\rangle$ (17)
@@
$=\frac{1}{\sqrt{d}^{d-1}}({\hat{a}}_{0}+{\hat{a}}_{1}+\cdots+{\hat{a}}_{d-1})^{l}({\hat{a}}_{0}+\omega{\hat{a}}_{1}+\cdots+\omega^{d-1-l}{\hat{a}}_{d-1})^{d-1-l}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle$
$=\frac{1}{\sqrt{d}^{d-1}}({\hat{a}}_{0}+{\hat{a}}_{1}+\cdots+{\hat{a}}_{d-1})^{l}(d-1-l)!\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}{\hat{a}}_{j_{1}}{\hat{a}}_{j_{2}}\cdots{\hat{a}}_{j_{d-1-l}}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle$
@@
Using
$\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}=(-1)^{d-1-l}\sum_{j=0}^{d-1}\omega^{(d-1-l)}=0,$ (19)
Eq. (B.2) becomes
@@
$=-\frac{l!}{\sqrt{d}^{d-1}}(d-1-l)\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}{\hat{a}^{\dagger}}_{j_{1}}$
$=(-1)^{d-1-l}\frac{l!(d-1-l)!}{\sqrt{d}^{d-1}}\sum_{j_{1}=0}^{d-1}\omega^{(d-1-l)j_{1}}{\hat{a}^{\dagger}}_{j_{1}}$

### v2

@@ 3. $X_{d}$: The $d$-level unitary generalization of $\sigma_{x}$ in qubit,
$\frac{1}{3^{9}}\Big{[}$ $6^{3}{\hat{a}^{\dagger}}_{00}{\hat{a}^{\dagger}}_{21}{\hat{a}^{\dagger}}_{12}{\hat{b}^{\dagger}}_{00}{\hat{b}^{\dagger}}_{21}{\hat{b}^{\dagger}}_{12}{\hat{c}^{\dagger}}_{00}{\hat{c}^{\dagger}}_{21}{\hat{c}^{\dagger}}_{12}$
$+6^{3}{\hat{a}^{\dagger}}_{20}{\hat{a}^{\dagger}}_{11}{\hat{a}^{\dagger}}_{02}{\hat{b}}^{\dagger}_{20}{\hat{b}}^{\dagger}_{11}{\hat{b}}^{\dagger}_{02}{\hat{c}^{\dagger}}_{20}{\hat{c}^{\dagger}}_{11}{\hat{c}^{\dagger}}_{02}$
$-(3\sqrt{3})^{3}\hat{X}^{\dagger}_{1}\hat{Y}^{\dagger}_{1}\hat{Z}^{\dagger}_{1}({\hat{a}^{\dagger}}_{1,1}{\hat{a}^{\dagger}}_{2,1}+{\hat{a}^{\dagger}}_{1,2}{\hat{a}^{\dagger}}_{2,0})$
×$~{}~{}~{}\times({\hat{b}^{\dagger}}_{1,1}{\hat{b}^{\dagger}}_{2,1}+{\hat{b}^{\dagger}}_{1,2}{\hat{b}^{\dagger}}_{2,0})({\hat{c}^{\dagger}}_{1,1}{\hat{c}^{\dagger}}_{2,1}+{\hat{c}^{\dagger}}_{1,2}{\hat{c}^{\dagger}}_{2,0})\Big{]}|vac\rangle$ (13)
After the postselection with $D_{F}$s and feedforwards, the final state that passes through $BS_{t}$ becomes
$\frac{6^{3}}{3^{9}}\Big{[}|0,0,0\rangle-3\sqrt{3}t^{3}|1,1,1\rangle+|2,2,2\rangle\Big{]}.$ (14)
Setting the transmissivity $t=\frac{1}{\sqrt{3}}$, the final state becomes the tripartite qutrit GHZ state. The success probability is given by
$P_{suc}(3,3)=3\times(\frac{2}{9})^{6}\sim 3.6\times 10^{-4}.$ (15)
@@
The linear optical logic gates in Fig. 1 can be implemented with any qudit encoding system of linear optics. Among them, we discuss the OAM and multi-rail encodings.
In the OAM encoding, it is straightforward to establish the scheme, because the OAM beam splitters [39], OAM-only Fourier transformation operators [40], and OAM high-dimensional $X$ gate [41] implement the $CN_{d}$, $F_{d}$, and $X_{d}$ gates respectively.
In the multi-rail encoding, a permutation of photon paths becomes the $CN_{d}$ gate and the $d$-partite port that Fourier transforms the spatial states of photons becomes the $F_{d}$ gate (see Appendix E). An advantage of the multirail encoding circuit is that $CN_{d}$ and $\overline{CN}_{d}$ is achieved with the permutation of paths, hence the final circuit consists of simpler optical elements, i.e., $d$-partite ports including BSs, permutations of paths, and photon detectors.
@@
Our scheme has several manifest advantages in using the qudit GHZ state as a resource for high-dimensional quantum information processing. First, as we have already demonstrated, our scheme enables more efficient generations of the resource states with the minimal number of single-photons. Given that 20 single photons can be successfully generated and controlled in linear optical systems [42], we expect relatively low-dimensional qudit GHZ states to be generated sooner or later using in laboratories or Quandela platform [43]. Second, our scheme can generate generalized qudit GHZ states, which is adequate to verify the optimal state for the violation of Bell-CHSH nonlocality test  [^1] While the maximally entangled state entails the maximal Bell-CHSH violation in $d=2$ systems, the scenario becomes more complicated for higher dimensional systems. Ref. [18, 19] shows that the maximal Bell-CHSH violation occurs for a non-maximally entangled state, which is also verified numerically for the $N=d=3$ case in Ref. [52]. Therefore, for the experiments of those maximal violations, schemes for generalized qudit GHZ states with relatively asymmetric amplitudes are indispensable. Our scheme easily achieve the condition by varying the transmissivity of beamspliters.. Third, our scheme consists of linear optical logic gates that any type of qubit encoding (e.g., optical angular momentum (OAM), multi-rail, time-bin, etc.) can realize.
It is worth emphasizing that the methodology of composing linear optical logic gates based on the linear quantum graph picture (LQG picture) [45, 37] (Appendix B.1 for a brief explanation) can be useful to find more general types of qudit multipartite entanglement. For example, Ref. [46] proposed heralded schemes for $N$-partite $N$-level symmetric and anti-symmetric states based on the method. We expect to construct more rigorous and general formalism of the LQG picture that will provide a more systematic path for the heralded generations of essential qudit entanglement such as absolutely maximally entangled states.
In addition, it will be an intriguing problem to find the upper bound of the success probability for generating $|GHZ_{N}^{d}\rangle$ with minimal photon number, $dN$. We expect that the mathematical technique to determine the required minimal photon number (Appendix D) could be exploited to search for the solution.
@@
SC is grateful to Prof. William J. Munro and Prof. Jung-Hoon Chun for their support on this research. SC is supported by National Research Foundation of Korea (NRF, 2019R1I1A1A01059964, RS-2023-00245747). JR acknowledges support from the National Research Foundation (NRF) of Korea grant funded by the Korea Government (Grant No. NRF-2022M3K2A1083890). YSK acknowledges support from the KIST institutional program (2E32941).
@@ Appendix A Operations of $CN_{d}$, the $d$-level CNOT gate between the spatial mode and the internal state
Unlike the conventional generalized CNOT gate that works on two different qudits, $CN_{d}$ and $\overline{CN}_{d}$ here transform the spatial mode of a given photon according to its internal state. They transform the two variables in the same way as the generalized CNOT gate between two qudits. The spatial mode and internal state play the role of the target and control states.
More explicitly, the operation of $CN_{d}$ given by Eq. (2) of the main content can be enumerated as the following table:
$|0\rangle_{0}\to|0\rangle_{0}$ $|1\rangle_{0}\to|1\rangle_{1}$cdots $|d-1\rangle_{0}\to|d-1\rangle_{d-1}$
$|0\rangle_{1}\to|0\rangle_{1}$ $|1\rangle_{1}\to|1\rangle_{2}$cdots $|d-1\rangle_{1}\to|d-1\rangle_{0}$
@@
$|0\rangle_{d-2}\to|0\rangle_{d-2}$ $|1\rangle_{d-2}\to|1\rangle_{d-1}$cdots $|d-1\rangle_{d-2}\to|d-1\rangle_{d-3}$
$|0\rangle_{d-1}\to|0\rangle_{d-1}$ $|1\rangle_{d-1}\to|1\rangle_{0}$cdots $|d-1\rangle_{d-1}\to|d-1\rangle_{d-2}$
@@ Appendix B Comprehensive analysis on the qudit $N$-partite GHZ state generation scheme
Here we provide a comprehensive explanation on the linear optical circuit for generating the qudit GHZ state given in Figure 2 of the main content. We first explain how we designed our scheme motivated by the graph picture of linear quantum networks (or linear quantum graph picture, LQG picture) [45, 37] and then show that the scheme indeed generates the expected target state.
B.1 From sculpting operators to photon subtraction operators
The LQG picture provides a set of correspondence relations between graphs and many-particle systems that linearly evolve. This graph picture was introduced to provide systematical methods to search for schemes that can generate genuine multipartite entanglement. Ref. [45] first introduced the LQG picture and found several postselected schemes for multipartite entangled states. Ref. [37] generalized the formalism and applied it to find schemes for generating multipartite entanglement with the sculpting scheme [47] (which generates an N -partite entangled state by applying tailored single-boson annihilation operators). For the qubit case, the graph solutions for generating entanglement with sculpting schemes in Ref. [37] can be implemented in linear optical systems probabilistically by a set of direct translation rule [38]. However, the sculpting scheme for generating high-dimensional GHZ state in Ref. [37] has remained conceptual without any suggestion to experimentally execute the high-dimensional sculpting process in many physical system. And this process for qudits is considerably more complicated than the qubit case in which the particle paths are simply split using polarized beam splitters (PBS) by the Hong-Ou-Mandel effect.
In this section, we explain the sculpting scheme for $|GHZ_{N}^{d}\rangle$ proposed in Ref. [37] and how to construct linear optical logic gates that execute the high-dimensional boson annihilation operations for generating the qudit GHZ state.
In Ref. [37], a boson sculpting scheme for $|GHZ_{N}^{d}\rangle$ is introduced as follows:
  1. 1.
    Prepare for $dN$ bosons in $N$ spatial modes so that each boson states are orthogonal to each other, e.g.,
    $|Init_{N,d}\rangle\equiv\prod_{j=1}^{N}({\hat{a}^{\dagger}}_{j,0}{\hat{a}^{\dagger}}_{j,1}\cdots{\hat{a}^{\dagger}}_{j,d-1})|vac\rangle,$ (16)
    where ${\hat{a}^{\dagger}}_{j,s}$ denotes a boson in the $j$th subsystem ($j\in\{1,2,\cdots,N\}$) with internal qudit state $s$ ($(\in\{0,1,\cdots,d-1\})$.
  2. 2.
    Apply a boson sculpting operator
    $\hat{A}_{N,d}=\Big{(}\frac{1}{\sqrt{2}}\Big{)}^{(d-1)N}({\hat{a}}_{1,{\tilde{0}}}-{\hat{a}}_{2,\widetilde{d-1}})^{d-1}({\hat{a}}_{2,{\tilde{0}}}-{\hat{a}}_{3,\widetilde{d-1}})^{d-1}({\hat{a}}_{3,{\tilde{0}}}-{\hat{a}}_{4,\widetilde{d-1}})^{d-1}\cdots({\hat{a}}_{N,{\tilde{0}}}-{\hat{a}}_{1,\widetilde{d-1}})^{d-1}.$ (17)
  3. 3.
    Using Identities 1 and 2 in Appendix Section B.2, the final state is given by the qudit GHZ state in the Fourier-transformed basis up to normalization:
    $\Big{(}\frac{1}{\sqrt{2}}\Big{)}^{(d-1)N}\Big{(}|\tilde{0},\tilde{0},\cdots,\tilde{0}\rangle+|\tilde{1},\tilde{1},\cdots,\tilde{1}\rangle+\cdots+|\widetilde{d-1},\widetilde{d-1},\cdots,\widetilde{d-1}\rangle\Big{)}$ (18)
The graph representation of the operator (17) is given by (see Table I of Ref. [37])
Decomposing the above graph, it consists of
[Uncaptioned image]$\end{gathered},~{}~{}~{}~{}(j\in\{1,2,\cdots,N\},~{}\textrm{$\oplus_{d}$ is addition mod $d$})$ (21)
which corresponds to the operator
$\frac{1}{\sqrt{2}^{d-1}}({\hat{a}}_{j,\tilde{0}}-{\hat{a}}_{j\oplus_{d}1,\widetilde{d-1}})^{d-1}.$ (22)
Therefore, we need a linear optical element that subtracts $(d-1)$ photons, whose photon state is the superposition of $|j,\tilde{0}\rangle$ and $|j\oplus_{d}1,\widetilde{d-1}\rangle$.
From the property of the $CN_{d}$ gate (Eq. (2) of the main content), we see that photons of state $\tilde{0}$ ($\widetilde{d-1}$) comes from the 1st ($(d-1)$-th) mode of a $CN_{d}$ gate. And we detect $(d-1)$ photons in the Fourier-transformed bases to herald the subtraction operator. By combining these observations, we can expect that the following correspondence holds from the LQG picture to the linear optical system:
Note that $X_{d}$ is attached to the $d$-th mode of the $CN_{d}$ gate from the $(j\oplus_{d}1)$-th mode to simplify the relative phases of the photons from the two different spatial modes. And photons from the $2$nd to the $(d-2)$-th modes of $CN_{d}$ gates must remain in the same subsystem.
@@ B.1 From sculpting operators to photon subtraction operators
By connecting the $N$ sets of the above linear optical element based on the structure of the bigraph (19), we obtain the scheme in Figure 2 of the main content.
@@
B.2 Demonstrations of the final state generated by the scheme in Figure 2 of the main content
@@
To demonstrate that the scheme in Figure 2 of the main content actually generates the qudit GHZ state, we need the following two identities that were first given in Ref. [37] without proofs. Due to their significance in calculating our circuit operation, we restate them here with proofs.
@@
$({\hat{a}}_{{\tilde{0}}})^{l}({\hat{a}}_{\widetilde{d-1}})^{d-1-l}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle=(-1)^{d-1-l}\frac{l!(d-1-l)!}{\sqrt{d}^{d-2}}{\hat{a}}^{\dagger}_{\widetilde{d-1-l}}|vac\rangle.$ (24)
Here all the operators are on the same subsystem, hence ${\hat{a}^{\dagger}}_{s}$ and ${\hat{a}}_{s}$ denote the creation and annihilation operators and $s$ is the internal qudit state of the operators.
@@
Using the following relations
${\hat{a}^{\dagger}}_{\tilde{j}}=\frac{1}{\sqrt{d}}({\hat{a}}_{0}+\omega^{j}{\hat{a}}_{1}+\cdots+\omega^{j(d-1)}{\hat{a}}_{d-1}),$
${\hat{a}}_{\tilde{j}}=\frac{1}{\sqrt{d}}({\hat{a}}_{0}+\omega^{-j}{\hat{a}}_{1}+\cdots+\omega^{-j(d-1)}{\hat{a}}_{d-1}),$ (25)
we have
@@
$=\frac{1}{\sqrt{d}^{d-1}}({\hat{a}}_{0}+{\hat{a}}_{1}+\cdots+{\hat{a}}_{d-1})^{l}({\hat{a}}_{0}+\omega{\hat{a}}_{1}+\cdots+\omega^{(d-1)}{\hat{a}}_{d-1})^{d-1-l}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle$
$=\frac{1}{\sqrt{d}^{d-1}}({\hat{a}}_{0}+{\hat{a}}_{1}+\cdots+{\hat{a}}_{d-1})^{l}\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{(j_{1}+j_{2}+\cdots+j_{d-1-l})}{\hat{a}}_{j_{1}}{\hat{a}}_{j_{2}}\cdots{\hat{a}}_{j_{d-1-l}}\prod_{s=0}^{d-1}{\hat{a}}^{\dagger}_{s}|vac\rangle$
@@
On the other hand, since
$\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}=0$ (27)
holds, Eq. (B.2) is simplified as
@@
$=-\frac{l!}{\sqrt{d}^{d-1}}(d-1-l)\sum_{\begin{subarray}{c}j_{1}\neq j_{2}\neq\cdots\\ \neq j_{d-1-l}\\ =0\end{subarray}}^{d-1}\omega^{j_{1}+j_{2}+\cdots+j_{d-1-l}}{\hat{a}^{\dagger}}_{j_{1}}|vac\rangle$
$=(-1)^{d-1-l}\frac{l!(d-1-l)!}{\sqrt{d}^{d-1}}\sum_{j_{1}=0}^{d-1}\omega^{(d-1-l)j_{1}}{\hat{a}^{\dagger}}_{j_{1}}|vac\rangle$